SOLVE
4x^2 -3x-7=0

Answer by rothauserc(4718)

4x^2 -3x-7 = 0
:
factor the polynomial
:
(4x-7) * (x+1) = 0
:
1) 4x - 7 = 0
:
4x = 7
:
x = 7/4
:
2) x +1 = 0
:
x = -1
:
*****************************************
x = 7/4 and x = -1
:
check our answers
:
1) x = 7/4
:
4*(7/4)^2 -3*(7/4) -7 = 0
:
4*49/16 -21/4 -7 = 0
:
196/16 -84/16 -7 = 0
:
112/16 -7 = 0
:
7 - 7 = 0
:
0 = 0
:
2) 4*(-1)^2 -3*(-1) -7 = 0
:
4 +3 -7 = 0
:
7 - 7 = 0
:
0 = 0
:
our answers check :-)
******************************************
: